OpenStudy (anonymous):

ok, so im doing cubed roots and stuff like that..i need help on this problem...(8x64) with an exponent 2/3.....the x meaning times

OpenStudy (anonymous):

I believe it would be 64

OpenStudy (anonymous):

thats what my book says, but can you explain a little of how to get that answer?

OpenStudy (anonymous):

(8 x 64)^(2/3) = ( 8 x 8 x 8)^(2/3) = (8^3)^ 2/3 3*2/3 =8^ = 8^2 = 64

OpenStudy (anonymous):

Is it clear now ????????

OpenStudy (anonymous):

yes, that was easier than what my teacher taught us. thank you!

OpenStudy (anonymous):

u r most welcome....☺